To help keep his barn warm on cold days, a farmer stores 805 kg of solar-heated water (Lf = 3.35E+5 J/kg) in barrels. For how many hours would a 2.12 kW electric space heater have to operate to provide the same amount of heat as the water does, when it cools from 10.1 to 0°C and completely freezes?
Answer #1

total heat required = mcdT + mL

= 805*(4186*10.1 + 3.35*10^5)

= 303709273 J

time required = Q/P

= 303709273 / 2120 = 143259.1 s

= 143259.0 / 3600 = 39.8 h

so it has to be operated for 39.8 h